# Free energy of the gas of spin 1/2 fermions beyond the second order and the Stoner phase transition

### v1

@@ 5. The particle-hole diagrams
analogous to (26) but with now different meaning of the symbols $\{{\mathbf{p}}\}$ and $\mathbf{p}$}]:
@@
$[{\mathbf{p}}]\equiv\varepsilon_{\mathbf{p}}-\tilde{\mu}_{-}^{(0)}-\varepsilon_{{\mathbf{q}}+{\mathbf{p}}}+\tilde{\mu}_{+}^{(0)}~\!,$
@@
(it is in fact the identity (35) but written with different symbols). Using the symmetry of this expression allows to write the expression for $F^{(3)ph}/V$ as the sum of two terms which in the zero temperature limit reproduce the two terms, $K_{1}$ and $K_{2}$, which in [14, 15] represented the contribution of the second diagram of Figure 2 to $E/V$.
@@
Making now the changes of the variables: first ${\mathbf{p}}_{1}+{\mathbf{q}}={\mathbf{k}}_{1}$, ${\mathbf{p}}_{1}=-{\mathbf{k}}_{2}$, and then in the $n_{+}$ terms of $\{{\mathbf{p}}_{2}\}$ the change ${\mathbf{p}}_{2}+{\mathbf{k}}_{1}+{\mathbf{k}}_{2}=-{\mathbf{p}}$ and ${\mathbf{p}}_{2}={\mathbf{p}}$ in the $n_{-}$ terms of $\{{\mathbf{p}}_{2}\}$, and then taking explicitly the integral over the cosine of the angle between $\mathbf{p}$} and ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}$ the expression for $F^{(3)ph}/V$ can be written in the form
@@
Again, the limits of the integrals in $K_{\delta}$ are determined by the condition that the zeroes of the arguments of the Dirac deltas treated as functions of the cosine of the angle between $\mathbf{p}$} and ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}$ lie between 1-1 and +1+1. And again the function $K_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})$ can be written down in a closed analytical form - see below.
@@
The expression (54) is finite although this is not immediately obvious: while the integrals defining the functions $K$ and $K_{\delta}$ are clearly (ultraviolet) finite owing to the presence of the distribution functions $n_{+}(p)$ and $n_{-}(p)$, the integral over $d^{3}{\mathbf{k}}_{1}$ has no such an exponentially suppressing factor. In addition, the presence of the factors $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$ in front of the functions $K$ and $K_{\delta}$ seem to imply the potential presence of a singularity in the integral over the directions of ${\mathbf{k}}_{1}$ (for $|{\mathbf{k}}_{1}|=|{\mathbf{k}}_{2}|$). To prove the finiteness of (54) we will first analyze the behavior of the difference $K^{2}-K_{\delta}^{2}/3$ in the limit $|{\mathbf{k}}_{1}|\rightarrow\infty$. To this end it is helpful to split the functions $K$ and $K_{\delta}$ into $K_{+}+K_{-}$ and $K_{\delta+}+K_{\delta-}$ (the subscripts $\pm$ corresponds to the $n_{\pm}$ distribution functions). The hardest to see is the convergence of the integral over $d^{3}{\mathbf{k}}_{1}$ involving the factor $K_{-}^{2}-K^{2}_{\delta-}/3$. The dangerous term $n_{-}({\mathbf{k}}_{2})[K_{-}^{2}-K^{2}_{\delta-}/3]$ comes from the term
@@
in (51). This, however is algebraically zero (just as the integrand of (36)) and, therefore, the factor $K_{-}^{2}-K^{2}_{\delta-}/3$ must be zero too. As any inaccuracy in the numerical evaluation of the integrals (55) and (56) could lead to a nonzero difference $K_{-}^{2}-K^{2}_{\delta-}/3$ and, therefore, to a (fake) nonconvergence of the integration over $d^{3}{\mathbf{k}}_{1}$, in the term with unity arising from $1-n_{+}({\mathbf{k}}_{1})]$ we simply replace $K^{2}-K^{2}_{\delta}/3$ by $K_{+}^{2}+2K_{+}K_{-}-(K_{\delta+}^{2}+2K_{\delta+}K_{\delta-})/3$. The rest of the terms are separately integrable in the limit $|{\mathbf{k}}_{1}|\rightarrow\infty$:
@@
because $\Delta^{2}_{1}$ grows like ${\mathbf{k}}_{1}^{2}$ as $|{\mathbf{k}}_{1}|\rightarrow\infty$. This secures the convergence of the integrals of the terms $K_{\delta+}^{2}$ and $2K_{\delta+}K_{\delta-}$. Similarly, it can be estimated that the integral over $p$ in $K_{+}$ behaves as $1/\Delta_{1}$ when $|{\mathbf{k}}_{1}|\rightarrow\infty$. Since each of the $K_{\pm}$ functions has the factor $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$ in front of it, the term $(K_{+})^{2}$ behaves for $|{\mathbf{k}}_{1}|\rightarrow\infty$ as $1/({\mathbf{k}}_{1}^{2}+{\mathbf{k}}_{1}\cdot{\mathbf{k}}_{2})^{2}$ and this secures the convergence of the integration over $d^{3}{\mathbf{k}}_{1}$. The term $2K_{+}K_{-}$, instead, behaves only as $1/({\mathbf{k}}_{1}^{2}+{\mathbf{k}}_{1}\!\cdot\!{\mathbf{k}}_{2})|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$, but the integration over the cosine of the angle between ${\mathbf{k}}_{1}$ and ${\mathbf{k}}_{2}$ kills the term of order $1/|{\mathbf{k}}_{1}|^{3}$ and the remaining terms are already integrable. This only power-like suppression of the integration of the term $K_{+}^{2}+2K_{+}K_{-}$ makes, however, numerical evaluation of the particle-hole contribution to the free energy more difficult and, therefore, potentially less accurate than the evaluation of the particle-particle one. As to the potentially singular factors $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$, one should first notice that the original expressions (48) and (49) as well as similar formulae giving the contributions of the $N$-th order particle-hole rings do not contain such singularities. They are due to the symmetrizations needed to arrive at the final formulae and must, therefore, cancel out (like the spurious singularities of the integrand in (36)) even if it is not directly evident. In the third order finitness of (54) can be seen as follows: since the integrals in the definitions of the $K_{\pm}$ and $K_{\delta\pm}$ functions are finite in the limit ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}\rightarrow{\mathbf{0}}$, the singularities of (54) have essentially the form $1/({\mathbf{k}}_{1}^{2}+{\mathbf{k}}_{2}^{2}+2|{\mathbf{k}}_{1}||{\mathbf{k}}_{2}|\xi)$ where $\xi$ is the cosine of the angle between ${\mathbf{k}}_{1}$ and ${\mathbf{k}}_{2}$. In the third order contribution considered here, after integration over $\xi$ they give rise to terms $\ln\!||{\mathbf{k}}_{1}|-|{\mathbf{k}}_{2}||$ which are integrable. [^5] In fact for $|{\mathbf{k}}_{1}|=|{\mathbf{k}}_{2}|$ the factors $\Delta_{1}$ and $\Delta_{2}$ behave as $\sqrt{1+\xi}$ and it can be checked (numerically) that the integrals in the functions $K_{+}$ and $K_{-}$ vanish then as $\sqrt{1+\xi}$, so these functions have therefore finite limits. The singularities reside only in the terms involving the functions $K_{\delta\pm}$. In the numerical evaluation of the integrals in (54) and in (60) it is however sufficient to impose a cutoff $|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|>\kappa$ and check that the results stabilize as $\kappa$ approaches zero. Thus the expression (54) is finite and we have checked that evaluated for $T=0$ (so that the Fermi distribution functions can be replaced by the step functions) it reproduces numerically in the entire range of polarizations $P$ the contribution to the ground state energy density of the third order particle-hole diagram of Figure 2 obtained in [14, 15].
@@
From this formal sum one has to subtract two first terms of the series: there is no order $C_{0}^{\rm R}$ particle-hole diagram and the order $(C_{0}^{\rm R})^{2}$ term is already (recall that the order $C_{0}^{2}$ contribution of the diagram shown in Figure 1 can be written either as a convolution of two $A$-blocks or of two $B$-blocks shown in the same Figure), in the properly renormalized form, included in (47). Thus the final form of the ressumed contributions of the particle-hole ring diagrams is
@@
One has to comment again on the finiteness of the expression (60). The singularities related to the factors $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$ are now harmless because the function $K$ is, as remarked, finite in the limit ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}\rightarrow{\mathbf{0}}$ and the singular function $K_{\delta}$ is now in the denominator and under the arctan function. As to the ultraviolet finiteness of (60), the integral of the factor $K$ explicitly subtracted in the square brackets is ultraviolet divergent being simply equivalent to the divergent expression (28). To ensure proper cancelation of this term in (60) for large $|{\mathbf{k}}_{1}|$ and/or $|{\mathbf{k}}_{2}|$ we expand the arctan function in $K_{\delta}$ and $K$ up to the sixth order (we have checked that taking more terms of the expansion does not change the result appreciably). The ultraviolet divergent terms linear in $K$ then disappear and, moreover, the expansion allows to implement the discussed trick with replacing $K^{2}-K_{\delta}^{2}/3$ by $K^{2}_{+}+2K_{+}K_{-}-(K^{2}_{\delta+}+2K_{\delta+}K_{\delta-})$.
In [15] we have compared the order $(k_{\rm F}a_{0})^{3}$ contributions to the ground state energy (i.e. to the free energy $F$ for zero temperature) of the particle-particle and of the particle-hole diagrams (shown in Figure 2) in the entire range of the polarization $P$ and found that the second one is not much smaller than the first one. In Figures 5, 6 and 7 we compare the magnitudes of the resummed contributions (47) and (60) to the free energy as functions of $k_{\rm F}a_{0}$ for three different temperatures and two representative values of the system’s polarization $P$. We also plot the sum of these two contributions. It again follows that for $k_{\rm F}a_{0}\sim 1$ the resummed contribution of the particle-hole diagrams is not much smaller than that of the particle-particle ones and is of the opposite sign. It is also clear that in this most important region the sum of the two contributions is significantly distorted compared to the resummed contribution of the particle-particle diagrams. This raises the question what impact the resummed contribution of the particle-hole ring diagrams has on the results reported in the papers [18, 20, 30].
@@
Figure 8: The difference $(F(P)-F(0))^{pp+ph}/N$ (in units $(3/5)\varepsilon_{\rm F}$ for $T=0$ and $T=0.2~\!T_{\rm F}$ as a function of the polarization $P=(N_{+}-N_{-})/N$ for different values of the gas parameter $k_{\rm F}a_{0}$.
Figures 8 and 9, analogous to Figures 3 and 4 illustrate the consequences of adding the resummed contribution (60) of the particle-hole ring diagrams to the free energy. The result is dramatic: the phase transition to the ordered state discussed in [18, 20, 30] completely disappears! Thus, even if the selection of the particle-particle ring diagrams as giving the dominant contribution can be (partially) justified by invoking the arguments, given in [35], based on using $1/2^{D/2}$ where $D$ is the number of space dimensions as the expansion parameter, they in practice do not turn out to be really dominant: the contribution of other subsets of diagrams (the number of such subsets beginning at a given order of the expansion grows with the order number) can, as our results show, change qualitatively the behavior of the thermodynamical potentials of the system of fermions close to the Feshbach resonance.
@@ 7. Conclusions
We have applied the systematic thermal (imaginary time) perturbative expansion to the effective (low energy) field theory to compute the free energy of the gas of interacting (nonrelativistic) spin $1/2$ fermions for arbitrary values of the gas polarization and temperatures not exceeding the Fermi temperature. We have shown how to circumvent the technical problem which previously prevented us from immediately extending such a computation beyond the second order in the gas parameter $k_{\rm F}a_{0}$ and have given explicit formulae for the order $(k_{\rm F}a_{0})^{3}$ contributions to the system’s free energy. It turned out that the analytical part of this computations is more transparent and easier than the corresponding direct computations of the ground state energy based on the formula (14) which gives only the zero temperature limit of the results obtained with the help of the thermal expansion. (Of course, numerical evaluation of the resulting expressions for a nonzero temperature is considerably more involved than for $T=0$).
@@
To obtain the complete order $(k_{\rm F}R)^{3}$ contribution to the free energy of the gas of fermions interacting through a truly repulsive spin-independent two-body potential (characterized by a length scale $R$) one would have to add the contributions arising from the operators of lower length dimension in the effective theory interaction term because such potentials naturally give the $p$-wave scattering length $a_{1}$ and the $s$-wave effective range $r_{0}$ comparable to the $s$-wave scattering length $a_{0}$. Instead of doing this, in this work we have profited from the simple structure of the contributions of the particle-particle and particle-hole ring diagrams and managed to give simple formulae for their contributions to the free energy resummed to all orders in the gas parameter $k_{\rm F}a_{0}$. These formulae apply therefore rather to cold gases of fermionic atoms (interacting through attractive poentials) close to the Feshbach resonance where their $s$-wave scattering length $a_{0}$ is made positive and much larger than the remaining scattering lengths and effective radii. Using these formulae we have first checked that including only the contributions of the particle-particle rings we reproduce for zero temperatures all the results obtained in [18]. In particular we confirm that in this approximation at zero temperature the transition to the ordered phase occurs for $k_{\rm F}a_{0}=0.79$ and that it is continuous. These results seem to agree well with the results of the dedicated Monte Carlo computations. Our formula would, however, allow to study also the thermal profile of the transition.
@@
However we have found that the phase transition to the ordered state completely disappears after including into the free energy the resummed contribution of the particle-hole ring diagrams - the minimum of the free energy is always for zero polarization. This may indicate at least that the agreement of the critical value of the gas parameter $k_{\rm F}a_{0}$ found in the papers [18, 20, 30] with its value obtained from the Monte Carlo simulations (done with attractive potentials tuned so that $a_{0}$ is positive and large) is just accidental. It may however also indicate that there is indeed no transition (in the metastable state) to the spin ordered phase if the true interaction is attractive (but $a_{0}$ is made positive and large by approaching a Feshbach resonance). The vanishing at zero temperature of the inverse spin susceptibility ($1/\chi\propto(\partial^{2}F/\partial P^{2})_{P=0}$) for some value of the gas parameter found in the Monte Carlo simulations (which in the case of attractive potentials are not as clean as in the repulsive case because of the need to exclude - in the finite volume, that is without taking the thermodynamic limit - an overlap with the true ground state of the considered system of atoms) [21] from which one infers its existence and its continuous character would then be misleading (the second order approximation to $F$ predicts analytically vanishing of the inverse spin susceptibility at $k_{\rm F}a_{0}=1.058$ whereas the transition is first order and occurs for $k_{\rm F}a_{0}=1.054$ [18]). If this hypothesis is true it would provide yet another (in addition to the formation of atomic dimers) reason for the failure to simulate the intinerant ferromagnetism with the help of cold atoms. In an case it is clear that more theoretical studies are needed to clarify the situation.

### v2

@@ 5. The particle-hole diagrams
analogous to (26) but now with a different meaning of the symbols $\{{\mathbf{p}}\}$ and $\mathbf{p}$}]:
@@
$[{\mathbf{p}}]\equiv\varepsilon_{\mathbf{p}}-\tilde{\mu}_{-}^{(0)}-\varepsilon_{{\mathbf{q}}+{\mathbf{p}}}+\tilde{\mu}_{+}^{(0)}~\!.$
@@
[it is in fact the identity (35) but written with different symbols]. Using the symmetry of this expression allows to write the expression for $F^{(3)ph}/V$ as the sum of two terms that in the zero temperature limit reproduce the two terms, $K_{1}$ and $K_{2}$, which in Refs. [14, 15] represented the contribution of the second diagram of Fig. 2 to $E/V$.
@@
Making now the changes of the variables: first ${\mathbf{p}}_{1}+{\mathbf{q}}={\mathbf{k}}_{1}$, and ${\mathbf{p}}_{1}=-{\mathbf{k}}_{2}$, and then in the $n_{+}$ terms of $\{{\mathbf{p}}_{2}\}$ the change ${\mathbf{p}}_{2}+{\mathbf{k}}_{1}+{\mathbf{k}}_{2}=-{\mathbf{p}}$ and ${\mathbf{p}}_{2}={\mathbf{p}}$ in the $n_{-}$ terms of $\{{\mathbf{p}}_{2}\}$, and then taking explicitly the integral over the cosine of the angle between $\mathbf{p}$} and ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}$ the expression for $F^{(3)ph}/V$ can be written in the form
@@
Again, the limits of the integrals in $K_{\delta}$ are determined by the condition that the zeros of the arguments of the Dirac deltas treated as functions of the cosine of the angle between $\mathbf{p}$} and ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}$ lie between 1-1 and +1+1. And again the function $K_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})$ can be written down in a closed analytical form - see below.
@@
The expression (54) is finite although this is not immediately obvious: while the integrals defining the functions $K$ and $K_{\delta}$ are clearly (ultraviolet) finite owing to the presence of the distribution functions $n_{+}(p)$ and $n_{-}(p)$, the integral over $d^{3}{\mathbf{k}}_{1}$ has no such exponentially suppressing factor. In addition, the presence of the factors $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$ in front of the functions $K$ and $K_{\delta}$ seem to imply the potential presence of a singularity in the integral over the directions of ${\mathbf{k}}_{1}$ (for $|{\mathbf{k}}_{1}|=|{\mathbf{k}}_{2}|$). To prove the finitness of (54) we will first analyze the behavior of the difference $K^{2}-K_{\delta}^{2}/3$ in the limit $|{\mathbf{k}}_{1}|\rightarrow\infty$. To this end it is helpful to split the functions $K$ and $K_{\delta}$ into $K_{+}+K_{-}$ and $K_{\delta+}+K_{\delta-}$ (the subscripts $\pm$ corresponds to the $n_{\pm}$ distribution functions). The hardest to see is the convergence of the integral over $d^{3}{\mathbf{k}}_{1}$ involving the factor $K_{-}^{2}-K^{2}_{\delta-}/3$. The dangerous term $n_{-}({\mathbf{k}}_{2})[K_{-}^{2}-K^{2}_{\delta-}/3]$ comes from the term
@@
in (51). This, however, is algebraically zero [just as the integrand of (36)] and, therefore, the factor $K_{-}^{2}-K^{2}_{\delta-}/3$ must be zero also. As any inaccuracy in the numerical evaluation of the integrals (55) and (56) could lead to a nonzero difference $K_{-}^{2}-K^{2}_{\delta-}/3$ and thus to a (fake) nonconvergence of the integration over $d^{3}{\mathbf{k}}_{1}$, in the term with unity arising from $1-n_{+}({\mathbf{k}}_{1})]$ we simply replace $K^{2}-K^{2}_{\delta}/3$ by $K_{+}^{2}+2K_{+}K_{-}-(K_{\delta+}^{2}+2K_{\delta+}K_{\delta-})/3$. The rest of the terms are separately integrable in the limit $|{\mathbf{k}}_{1}|\rightarrow\infty$:
@@
because $\Delta^{2}_{1}$ grows like ${\mathbf{k}}_{1}^{2}$ as $|{\mathbf{k}}_{1}|\rightarrow\infty$. This secures the convergence of the integrals of the terms $K_{\delta+}^{2}$ and $2K_{\delta+}K_{\delta-}$. Similarly, it can be estimated that the integral over $p$ in $K_{+}$ behaves as $1/\Delta_{1}$ when $|{\mathbf{k}}_{1}|\rightarrow\infty$. Since each of the $K_{\pm}$ functions has the factor $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$ in front of it, the term $(K_{+})^{2}$ behaves for $|{\mathbf{k}}_{1}|\rightarrow\infty$ as $1/({\mathbf{k}}_{1}^{2}+{\mathbf{k}}_{1}\cdot{\mathbf{k}}_{2})^{2}$ and this secures the convergence of the integration over $d^{3}{\mathbf{k}}_{1}$. The term $2K_{+}K_{-}$, instead, behaves only as $1/({\mathbf{k}}_{1}^{2}+{\mathbf{k}}_{1}\!\cdot\!{\mathbf{k}}_{2})|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$, but the integration over the cosine of the angle between ${\mathbf{k}}_{1}$ and ${\mathbf{k}}_{2}$ kills the term of order $1/|{\mathbf{k}}_{1}|^{3}$ and the remaining terms are already integrable. However, this only power-like suppression of the integration of the term $K_{+}^{2}+2K_{+}K_{-}$ makes numerical evaluation of the particle-hole contribution to the free energy more difficult and therefore potentially less accurate than the evaluation of the particle-particle one. As to the potentially singular factors $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$, one should first notice that the original expressions (48) and (49) as well as similar formulas giving the contributions of the $N$-th order particle-hole rings do not contain such singularities. They are due to the symmetrizations needed to arrive at the final formulas and must therefore cancel out [like the spurious singularities of the integrand in (36)] even if it is not directly evident. In the third order the finiteness of (54) can be seen as follows: since the integrals in the definitions of the $K_{\pm}$ and $K_{\delta\pm}$ functions are finite in the limit ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}\rightarrow{\mathbf{0}}$, the singularities of (54) have essentially the form $1/({\mathbf{k}}_{1}^{2}+{\mathbf{k}}_{2}^{2}+2|{\mathbf{k}}_{1}||{\mathbf{k}}_{2}|\xi)$ where $\xi$ is the cosine of the angle between ${\mathbf{k}}_{1}$ and ${\mathbf{k}}_{2}$. In the third order contribution considered here, after integration over $\xi$ they give rise to terms $\ln\!||{\mathbf{k}}_{1}|-|{\mathbf{k}}_{2}||$ which are integrable. [^5] In fact for $|{\mathbf{k}}_{1}|=|{\mathbf{k}}_{2}|$ the factors $\Delta_{1}$ and $\Delta_{2}$ behave as $\sqrt{1+\xi}$ and it can be checked (numerically) that the integrals in the functions $K_{+}$ and $K_{-}$ vanish then as $\sqrt{1+\xi}$, so these functions have finite limits. The singularities reside only in the terms involving the functions $K_{\delta\pm}$. In the numerical evaluation of the integrals in (54) and (60) it is however sufficient to impose a cutoff $|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|>\kappa$ and check that the results stabilize as $\kappa$ approaches zero. Thus the expression (54) is finite and we have checked that evaluated for $T=0$ (so that the Fermi distribution functions can be replaced by the step functions) it reproduces numerically in the entire range of polarizations $P$ the contribution to the ground state energy density of the third order particle-hole diagram of Figure 2 obtained in [14, 15].
@@
From this sum one has to subtract two first terms of the series: there is no order $C_{0}^{\rm R}$ particle-hole diagram and the order $(C_{0}^{\rm R})^{2}$ term is already (recall that the order $C_{0}^{2}$ contribution of the diagram shown in Fig. 1 can be written either as a convolution of two $A$-blocks or of two $B$-blocks shown in the same figure), in the properly renormalized form, included in (47). Thus the final form of the resummed contributions of the particle-hole ring diagrams is
@@
One has to comment again on the finitness of the expression (60). The singularities related to the factors $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|$ are now harmless because the function $K$ is, as remarked, finite in the limit ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}\rightarrow{\mathbf{0}}$ and the singular function $K_{\delta}$ is now in the denominator and under the arctan function. As to the ultraviolet finiteness of (60), the integral of the factor $K$ explicitly subtracted in the square brackets is ultraviolet divergent being simply equivalent to the divergent expression (28). To ensure proper cancellation of this term in (60) for large $|{\mathbf{k}}_{1}|$ and/or $|{\mathbf{k}}_{2}|$ we expand the arctan function in $K_{\delta}$ and $K$ up to the sixth order (we have checked that taking more terms of the expansion does not change the result appreciably). The ultraviolet divergent terms linear in $K$ then disappear and, moreover, the expansion allows to implement the discussed trick with replacing $K^{2}-K_{\delta}^{2}/3$ by $K^{2}_{+}+2K_{+}K_{-}-(K^{2}_{\delta+}+2K_{\delta+}K_{\delta-})$.
In Ref. [15] we have compared the order $(k_{\rm F}a_{0})^{3}$ contributions to the ground state energy (i.e. to the free energy $F$ for zero temperature) of the particle-particle and of the particle-hole diagrams (shown in Fig. 2) in the entire range of the polarization $P$ and found that the second one is not much smaller than the first one. In Figs. 5-7 we compare the magnitudes of the resummed contributions (47) and (60) to the free energy as functions of $k_{\rm F}a_{0}$ for three different temperatures and two representative values of the system’s polarization $P$. We also plot the sum of these two contributions. It again follows that for $k_{\rm F}a_{0}\sim 1$ the resummed contribution of the particle-hole diagrams is not much smaller than that of the particle-particle ones and is of the opposite sign. It is also clear that in this most important region the sum of the two contributions is significantly distorted compared to the resummed contribution of the particle-particle diagrams. This raises the question of what impact the resummed contribution of the particle-hole ring diagrams has on the results reported in the Refs. [18, 19, 30].
@@
Figure 8: The difference $(F(P)-F(0))^{pp+ph}/N$ (in units $(3/5)\varepsilon_{\rm F}$ for $T=0$ and $T=0.2~\!T_{\rm F}$ as a function of the polarization $P=(N_{+}-N_{-})/N$ for different values (indicated in the panels) of the gas parameter $k_{\rm F}a_{0}$.
Figures 8 and 9, analogous to Figures 3 and 4 illustrate the consequences of adding the resummed contribution (60) of the particle-hole ring diagrams to the free energy. The result is dramatic: the phase transition to the ordered state discussed in Refs. [18, 19, 30] completely disappears! Thus, even if the selection of the particle-particle ring diagrams as giving the dominant contribution can be (partially) justified by invoking the arguments, given in Ref. [35], based on using $1/2^{D/2}$ where $D$ is the number of space dimensions, as the expansion parameter, they in practice do not turn out to be really dominant: the contribution of other subsets of diagrams (the number of such subsets beginning at a given order of the expansion grows with the order number) can, as our results show, change qualitatively the behavior of the thermodynamic potentials of the system of fermions close to the Feshbach resonance.
@@ 7. Conclusions
We have applied the systematic thermal (imaginary time) perturbative expansion to the effective (low energy) field theory to compute the free energy of the gas of interacting (nonrelativistic) spin $1/2$ fermions for arbitrary values of the gas polarization and temperatures not exceeding the Fermi temperature. We have shown how to circumvent the technical problem which previously prevented us from immediately extending such a computation beyond the second order in the gas parameter $k_{\rm F}a_{0}$ and have given explicit formulas for the order $(k_{\rm F}a_{0})^{3}$ contributions to the system’s free energy. It turned out that the analytical part of this computation is more transparent and easier than the corresponding direct computation of the ground state energy based on the formula (14) which gives only the zero temperature limit of the results obtained with the help of the thermal expansion. (Of course, numerical evaluation of the resulting expressions for a nonzero temperature is considerably more involved than for $T=0$).
@@
To obtain the complete order $(k_{\rm F}R)^{3}$ contribution to the free energy of the gas of fermions interacting through a truly repulsive spin-independent two-body potential (characterized by a length scale $R$) one would have to add the contributions arising from the operators of lower length dimension in the effective theory interaction term because such potentials naturally give the $p$-wave scattering length $a_{1}$ and the $s$-wave effective range $r_{0}$ comparable to the $s$-wave scattering length $a_{0}$. Instead of doing this, in this work we have profited from the simple structure of the contributions of the particle-particle and particle-hole ring diagrams and managed to give simple formulas for their contributions to the free energy resummed to all orders in the gas parameter $k_{\rm F}a_{0}$. The obtained formulas therefore apply rather to cold gases of fermionic atoms (interacting through attractive potentials) close to the Feshbach resonance where their $s$-wave scattering length $a_{0}$ is made positive and much larger than the remaining scattering lengths and effective radii. Using these formulas we have first checked that including only the contributions of the particle-particle rings we reproduce for zero temperatures all the results obtained in Ref. [18]. In particular we confirm that in this approximation at zero temperature the transition to the ordered phase occurs for $k_{\rm F}a_{0}=0.79$ and that it is continuous. These results seem to agree well with the results of the dedicated Monte Carlo computations. Our formula would allow to study also the thermal profile of the transition, if it really existed.
@@
However we have found that the phase transition to the ordered state completely disappears after including into the free energy the resummed contribution of the particle-hole ring diagrams - the minimum of the free energy is always for zero polarization. This may at least indicate that the agreement of the critical value of the gas parameter $k_{\rm F}a_{0}$ found in Refs. [18, 19, 30] with its value obtained from the Monte Carlo simulations (done with attractive potentials tuned so that $a_{0}$ is positive and large) is just accidental. It may however also indicate that there is indeed no transition (in the metastable state) to the magnetically ordered phase if the true interaction is attractive (but $a_{0}$ is made positive and large by approaching a Feshbach resonance). The vanishing at zero temperature of the inverse spin susceptibility ($1/\chi\propto(\partial^{2}F/\partial P^{2})_{P=0}$) for some value of the gas parameter found in the Monte Carlo simulations (which in the case of attractive potentials are not as clean as in the repulsive case because of the need to exclude - in the finite volume and a finite number of fermions, that is without taking the thermodynamic limit - an overlap with the true ground state of the considered system of atoms) [21] from which one infers its existence and its continuous character would then be misleading (the second order approximation to $F$ predicts analytically vanishing of the inverse spin susceptibility at $k_{\rm F}a_{0}=1.058$ whereas the transition is first order and occurs for $k_{\rm F}a_{0}=1.054$ [18]). If this hypothesis is true it would provide yet another (in addition to the formation of atomic dimers) reason for the failure to simulate the itinerant ferromagnetism with the help of cold atoms. In an case it is clear that more theoretical studies are needed to clarify the situation. An obvious step in this direction would be extending the computation of ref. [10] of the order $(k_{\rm F}a_{0})^{4}$ contribution to nonequal densities of spin up and spin down fermions.
Data Availability Statement
The source code for this project is publicly available on Zenodo [36].
Acknowledgements
This research was funded, in whole or in part, by l’Agence Nationale de la Recherche (ANR), project ANR-23-CE31-0019..